OpenStudy (anonymous):

Why am i getting the wrong answer x^11y^10/x^-3y^-1 I get x^14, y^10 a -mn (quotient power i think).. so 11 - (-3) = 14 and 10 - (-1) = 11.. unless for some reason the M and N are swapped with the y and i have to do -1 + 10 or something, please explain

OpenStudy (michele_laino):

please use this rule: \[\frac{ y ^{n} }{ y ^{-m} }=y ^{n-(-m))}=y ^{n+m}\]

OpenStudy (anonymous):

@Michele_Laino i am 10 - (-1) is 11 isn't it?

OpenStudy (michele_laino):

yes, it is

OpenStudy (jhannybean):

\[\frac{x^{11}y^{10}}{x^{-3}y^{-1}}\]\[\frac{x^{11}y^{10}}{\dfrac{1}{x^3}\dfrac{1}{y}} = (x^{11}x^3)(y^{10}y^1)\]and then remember....\(x^m \cdot x^n = x^{m+n}\)

OpenStudy (anonymous):

@Michele_Laino then i am confused why all the calculators i use to test my answer gives me the +9 instead +11

OpenStudy (michele_laino):

I thin becaus your expression can be write as @Jhannybean wrote

OpenStudy (michele_laino):

oops... I think because...

OpenStudy (anonymous):

so am i right? the answer should be (x^14)(y^11) once simiplifed?

OpenStudy (jhannybean):

9 does not make any sense whatsoever.

OpenStudy (jhannybean):

Yes, you are right.

OpenStudy (michele_laino):

yes I think!

OpenStudy (anonymous):

wolffram gives me wrong answer then :s, thanks!!

OpenStudy (michele_laino):

of course if your expression is: \[x ^{11}\frac{ y ^{10} }{ x ^{3} }y ^{-1}\] then your answer is another

OpenStudy (jhannybean):

Then you would have :\[\frac{x^{11}}{1}\cdot \frac{y^{10}}{x^3} \cdot \frac{1}{y} = x^8y^9\]